# Joint Attention Mechanism Learning to Facilitate Opto-physiological Monitoring during Physical Activity

### v1

@@ 4.2 Experimental Results
Ablation study: The ablation study is performed to evaluate the performance of AM-GAN using the in-house LU dataset. The following ablation study is covered: (I) without triaxial ACC as input; (II) without discriminator; (III) without attention mechanism; and (IV) the whole AM-GAN method. The results are presented in Table 4
Attention Mechanism Discriminator triaxial ACC Input $\overline{Err1}$(HR) $\overline{Err2}$(HR) $\overline{Err1}$(RR) $\overline{Err2}$(RR)
w/o w/o w/o 3.81 3.90 4.99 14.86
w/o w. w/o 3.22 3.69 4.63 14.43
w/o w/o w. 2.99 3.03 4.11 13.98
w/o w. w. 2.45 2.71 3.65 13.12
w. w/o w. 1.89 1.73 3.01 12.35
w. w. w. 1.37 1.29 2.49 10.78
@@
Table 4: Ablation study of proposed AM-GAN in terms of attention mechanism, discriminator and triaxial ACC input for HR and RR validation on in-house LU dataset.
@@
Comparing the results by (I) and (II), we can observe that using triaxial ACC input can reduce the $\overline{Err1}$($\overline{Err2}$) by 0.23(0.66) beats/min(%) for HR validation and 0.52(0.45) breaths/min(%) for RR. This suggests that the triaxial ACC input as multi-sensor fusion is helpful for improving HR and RR validation accuracy. Similarly, using the attention mechanism for multi-sensor fusion can greatly improve HR and RR validation accuracy. These results indicate that the attention mechanism integrates inputs from multiple sensors for multi-sensor feature fusion to learn a good distribution about the MAs mixed with the PPG signals. It leads to better PPG feature disentanglement.
@@
Figure 10: Raw signal and de-noised PPG signals in time and frequency domains. (a) Time-domain raw PPG signal. (b) Frequency-domain raw PPG signal (P) and triaxial ACC signals (Ax, Ay and Az). (c,d) Time and frequency-domain generated de-noised PPG signal from PPG-GAN [23]. (e,f) Time and frequency-domain de-noised PPG signal from MR [15]. (g,h) Time and frequency-domain generated de-noised PPG signal from AM-GAN.
@@
Additionally we present a visualisation of the synthetic outcomes by comparing results (III) and (IV). The generated de-noised PPG signals are demonstrated in time and frequency domains by the without attention mechanism method (PPG-GAN), the SOTA MA-removal algorithm (MR) and the proposed AM-GAN in Fig. 10. This demonstration presents a complex example from the MW dataset in which the period of the HR and the motion period detected by the three-axis accelerometer sensor align closely, as depicted in Fig. 10(b). The MR method, which employs triaxial ACC as the motion reference, inadvertently filters out the real HR frequency owing to the exact overlap between the motion reference period and the HR period, as illustrated in Fig. 10(e) and (f). PPG-GAN is not effective in dealing with this scenario, as shown in Fig. 10(c) and (d). This also explains why the AM-GAN is able to work as a noise-reduced model across varying intensities of exercise to enhance the robustness in calculating physiological parameters, i.e., HR and RR. Additionally, the loss incurred by the generator in both PPG-GAN and AM-GAN during the training and validation phases for the LU dataset is depicted in Fig. 11. Notably, both models employ an identical loss function for their generators, (7). The proposed AM-GAN exhibits lower loss values and a superior alignment of the training loss curve with the validation loss curve, indicating a more effective model generalisation. Conversely, the PPG-GAN without the attention mechanism demonstrates a minor degree of overfitting, as evidenced by the analysis in Fig. 11(a).
Figure 11: Generator loss comparison of (III) and (IV). (a)PPG-GAN. (b) AM-GAN.
@@ 5 Discussion
The study has created an adversarial learning strategy with an attention mechanism to obtain clinically acceptable physiological parameters through PPG signals acquired from those subjects engaged in exercise at varying intensities. In these scenarios of high intensity, the measured signals exhibit substantial distortion, primarily due to the incorporation of MAs, a situation exacerbated by the potential overlap between motion frequencies and those found within the cardiac waveform. It is imperative, therefore, that MAs are eliminated prior to the assessment of any physiological parameter. The output from the triaxial accelerometer to act as motion reference signals was acquired along with four-channel PPG signals of mOEPS, and MAs were removed using the adversarial network with an attention mechanism (AM-GAN). The values for HR and RR were extracted from the generated MA-free signal showing the AM-GAN to have accurate vital signs even in the state of higher-intensity exercise. The analysis of these outcomes across three distinct datasets reveals a substantial enhancement in accuracy attributable to the consolidation of the modified GAN and the multi-head cross-attention mechanism.
The quality of the pulsatile waveform, quantified by the Pearson correlation coefficient $R$, is illustrated in Table 5. Fig. 5(a) and (b) evidence a high degree of consistency between the generated PPG signals and the reference signal across all LU datasets ($\overline{\langle R\rangle}>90$). The incorporation of the in-house LU dataset has facilitated the acquisition of benchmark readings for HR and RR. This benchmarking is crucial for assessing the accuracy of the AM-GAN method in extracting these vital measurements. The IEEE-SPC dataset and PPGDalia dataset served as a benchmark for comparing the performance of the AM-GAN against SOTA algorithms.
@@
It is presented that the AM-GAN has robust generalisation capabilities in scenarios characterised by unknown noise and MAs. However, some of the subject data have posed challenges with high $Err1$ values, such as S5, S6 and S8. A comparison of S5, S6 and S8 demonstrates that the lower signal quality of collected data and out-of-distribution HR values could cause the high $Err1$ values. Therefore, experimental outcomes using the in-house LU dataset, PPGDalia dataset, and IEEE-SPC dataset indicate that the proposed AM-GAN effectively eliminates MAs and improves the quality of noisy input PPG signals. The resulting clean PPG signal enables the determination of HR and RR with an accuracy exceeding 95% when compared to gold standard references. Additionally, the ablation study is performed to validate the interpretability of the AM-GAN method, as shown in Table 4. The visualisation of the generated de-noised PPG signals also presents better performance in the challenge scenarios, i.e., high-intensity exercise.
Although we achieved highly satisfactory results, there is still a lot of room to be improved. Firstly, we mainly focus on removing motion artefacts due to the availability of these published datasets. However, other artefacts, like muscular artefacts, may need to be eliminated in some scenarios. Second, we need the reference PPG signal from the SOTA MA-removal algorithms during the training phase. The performance of AM-GAN could significantly drop since it is bound to be inferior to generate the ground truth. There could be a rare straightforward solution to deliver quality ground truth yet its impact could be mitigated. The usage of unique training techniques, model architectures and model layers could decrease the data requirements in the training process.
@@ 6 Conclusion
The cardiac signal plays a crucial role in the computation of physiological parameters, including HR and RR. The study presents an AM-GAN approach to extract high-quality pulsatile waveforms for opto-physiological monitoring. The AM-GAN framework combines the generative adversarial and attention mechanisms, with error losses defined in both the time and frequency domains.
A noisy PPG signal, together with the triaxial ACC signals, comprises the network inputs; its output is an enhanced pulsatile waveform through the deep generative model that can be easily incorporated with existing opto-physiological monitoring methods.
The experimental results on two in-house datasets and one public dataset demonstrate that the AM-GAN consistently improves the quality of noisy input PPG signals to generate quality pulsatile waveforms, thus achieving accurate physiological parameters, i.e., HR and RR, both within-dataset and cross-dataset. The comparison results with other recently reported PPG methods, such as $Err1(Err2)$, verify the better performance of AM-GAN to calculate the accuracy of HR by generating high-quality pulsatile waveforms.
The AM-GAN has demonstrated its ability to learn an MA removal algorithm (here the MR algorithm, [15]), providing a fixed structure that can be updated as more effective algorithms are developed. Interestingly, it has also shown greater capability than the original MR algorithm in handling more complex situations, such as instances where cardiac cycle frequency coincides with the motion frequency during a higher-intensity exercise. While the results presented in the study are relatively preliminary, these efforts are significant in broadening the application spectrum of opto-physiological monitoring.
@@ Acknowledgments
The authors would like to acknowledge the support of Loughborough University in the conduction of this study and the shared datasets from the IEEE-SPC and PPGDalia.

### v2

@@ 4.2 Experimental Results
Ablation study: The ablation study is performed to evaluate the performance of AM-GAN using the in-house LU dataset. The following ablation study is covered: (I) without triaxial ACC as input; (II) without discriminator; (III) without attention mechanism; and (IV) the whole AM-GAN method. The results are presented in Table 5.
@@
Table 5: Ablation study of proposed AM-GAN in terms of attention mechanism, discriminator and triaxial ACC input for HR and RR validation on in-house LU dataset and SpO2 on in-house C2 dataset.
Attention Mechanism Discriminator triaxial ACC Input $\overline{Err1}$(HR) $\overline{Err2}$(HR) $\overline{Err1}$($R$) $\overline{Err2}$($R$) $\overline{Err1}$(SpO2) $\overline{Err2}$(SpO2)
3.81 3.90 4.99 14.86 3.67 3.58
3.22 3.69 4.63 14.43 3.11 3.21
2.99 3.03 4.11 13.98 2.86 2.99
2.45 2.71 3.65 13.12 2.73 2.80
1.89 1.73 3.01 12.35 2.05 1.97
1.37 1.29 2.49 10.78 1.65 1.79
@@
Comparing the results of (I) and (II), we can observe that using triaxial ACC input can reduce the $\overline{Err1}$($\overline{Err2}$) by 0.23(0.66) beats/min(%) for HR validation, 0.52(0.45) breaths/min(%) for RR and 0.25(0.22) % for SpO2. This suggests that the triaxial ACC input as multi-sensor fusion is helpful for improving HR, RR, and SpO2 validation accuracy. Similarly, using the attention mechanism for multi-sensor fusion can greatly improve HR, RR, and SpO2 validation accuracy. These results indicate that the attention mechanism integrates inputs from multiple sensors for multi-sensor feature fusion to learn a good distribution of the MAs mixed with the PPG signals. It leads to better PPG feature disentanglement.
@@
Figure 11: Raw signal and de-noised PPG signals in time and frequency domains. (a) Time-domain raw PPG signal. (b) Frequency-domain raw PPG signal (P) and triaxial ACC signals (Ax, Ay and Az). (c,d) Time and frequency-domain generated de-noised PPG signal from PPG-GAN [b29zheng2022ppg]. (e,f) Time and frequency-domain de-noised PPG signal from MR [b21zheng2023rapid]. (g,h) Time and frequency-domain generated de-noised PPG signal from AM-GAN.
@@
Additionally, we present a visualisation of the synthetic outcomes by comparing results (III) and (IV). The generated de-noised PPG signals are demonstrated in time and frequency domains by the without attention mechanism method (PPG-GAN), the SOTA MA-removal algorithm (MR) and the proposed AM-GAN in Fig. 11. This demonstration presents a complex example from the MW dataset in which the period of the HR and the motion period detected by the three-axis accelerometer sensor align closely, as depicted in Fig. 11(b). The MR method, which employs triaxial ACC as the motion reference, inadvertently filters out the real HR frequency owing to the exact overlap between the motion reference period and the HR period, as illustrated in Fig. 11(e) and (f). PPG-GAN is not effective in dealing with this scenario, as shown in Fig. 11(c) and (d). This also explains why the AM-GAN is able to work as a noise-reduced model across varying intensities of exercise to enhance the robustness in calculating physiological parameters, i.e., HR and RR. As presented in Fig. 12, the attention heatmap provides interpretability regarding which signal segments contribute the most to the MA-removal process. The attention weights learned by the model specifically reflect the relative importance assigned to different segments of the input raw PPG signal. Also, the red colouration in Fig. 12(a) denotes regions where the model places higher attention, particularly focusing on MA segments. To evaluate the sensitivity of the PPG-GAN to the choice of optimizer and learning rate (LR), we also conducted an ablation study (Table 6) in which the model was trained using three widely adopted optimizer, i.e., SGD (with momentum set to 0.9), AdamW, and Adam, each tested at three different learning rates: 0.01, 0.001, and 0.0002. All experiments were performed using the IEEE-SPC dataset.
Figure 12: Visualisation of attention maps is presented where the brighter parts indicate regions to which the generator pays more attention. (a) The attention map overlays highlight the raw PPG signal segments. (b) Reconstructed PPG signal from AM-GAN.
Table 6: Comparison of different optimizers and learning rates.
[…]
AdamW 0.01 $2±\pm2$.15
[…]
Adam (ours) 0.0002 $\pm1.81$
@@ 5 Discussion
The study has researched and developed an adversarial learning strategy with an attention mechanism to obtain clinically acceptable physiological parameters through PPG signals acquired from those subjects engaged in exercise at varying intensities. In high-intensity scenarios, the measured signals exhibit significant distortion, primarily due to motion artefacts (MAs), which are further exacerbated by the potential overlap between motion frequencies and cardiac waveform frequencies. It is essential to eliminate MAs before evaluating any physiological parameters. The output from the triaxial accelerometer to act as motion reference signals was acquired along with four-channel PPG signals of mOEPS, and MAs were removed using the adversarial network with an attention mechanism (AM-GAN). The values for HR, RR, and SpO2 were extracted from the generated MA-free signal, showing the AM-GAN to have accurate vital signs even in the state of higher-intensity exercise. The analysis of these outcomes across three distinct datasets reveals a substantial enhancement in accuracy attributable to the consolidation of the modified GAN and the multi-head cross-attention mechanism.
The pulsatile waveform, quantified by the Pearson correlation coefficient $R$, is illustrated in Table 5. Fig. 5(a) and (b) evidence a high degree of consistency between the generated PPG signals and the reference signal across all LU datasets ($\overline{\langle R\rangle}>90$). The incorporation of the in-house LU dataset has facilitated the acquisition of benchmark readings for HR and RR. This benchmarking is crucial for assessing the accuracy of the AM-GAN method in extracting these vital measurements. The IEEE-SPC dataset and PPGDalia dataset served as a benchmark for comparing the performance of the AM-GAN against SOTA algorithms.
@@
AM-GAN also has a better performance for cross-dataset (PPGDalia) HR validation with $\overline{Err1}$ of $3.86$ and $SD$ of $2.02$ beats/min compared with the SOTA methods, as shown in Table 3. It is presented that the AM-GAN has robust generalisation capabilities in scenarios characterised by unknown noise and MAs. However, some of the subject data have posed challenges with high $Err1$ values, such as S5, S6 and S8. A comparison of S5, S6 and S8 demonstrates that the lower signal quality of collected data and out-of-distribution HR values could cause the high $Err1$ values. Therefore, experimental outcomes using the in-house LU dataset, C2 dataset, PPGDalia dataset, and IEEE-SPC dataset indicate that the proposed AM-GAN effectively eliminates MAs and improves the quality of noisy input PPG signals. The resulting clean PPG signal enables the determination of HR, RR and SpO2 with an accuracy exceeding 95% when compared to gold standard references. Additionally, the ablation study is performed to validate the interpretability of the AM-GAN method, as shown in Table 5. The visualisation of the generated de-noised PPG signals also presents better performance in the challenge scenarios, i.e., high-intensity exercise.
Although we achieved highly satisfactory results, there is still a lot of room for improvement. First, our work primarily focuses on removing motion artefacts due to the availability of such published datasets. However, other artefacts, such as muscular artefacts, may also need to be addressed in certain scenarios. Second, our approach requires a reference PPG signal from state-of-the-art MA removal algorithms during the training phase. The performance of AM-GAN may be significantly affected if the generated ground truth is of suboptimal quality. While getting high-quality ground truth signals can be challenging, their impact may be mitigated. Employing advanced training techniques, innovative model architectures, and specialized model layers could reduce the dependence on large amounts of training data.
In the current state, the AM-GAN generator has a significant number of parameters, leading to a memory footprint of approximately 30 MB and requiring around 2 GFLOPs per inference. On a standard CPU, this physiological monitoring results (HR, RR and SpO2) in an inference latency of 500ms. For further deployment and optimisation, first, we would explore model pruning to remove redundant weights and channels from the generator. Techniques like magnitude-based pruning could significantly reduce the model’s size and computational load. Next, we would apply post-training quantisation. By converting the model’s weights from 32-bit floating-point numbers to 8-bit integers (INT8), we can achieve an almost $4\times$ reduction in model size and a significant speed-up in inference. Furthermore, we could employ knowledge distillation. Here, our AM-GAN would act as a ’teacher’ model. We would train a much smaller, lightweight ’student’ generator to mimic the output of the teacher. Finally, the optimised model would be compiled using a framework like TensorFlow Lite or PyTorch Mobile. These tools are specifically designed to run models efficiently on edge devices, taking advantage of hardware accelerators like Neural Processing Units (NPUs), which are becoming more common in wearable devices.
@@ 6 Conclusion
The cardiac signal is critical for computing physiological parameters such as HR, RR, and SpO2. This study introduces the AM-GAN framework, which integrates generative adversarial networks with attention mechanisms (AM-GAN), specifically designed to extract high-quality pulsatile waveforms for enhanced opto-physiological monitoring. The AM-GAN leverages both time- and frequency-domain error losses to effectively enhance noisy PPG signals, using these signals and associated triaxial ACC data as input.
Experimental evaluations on two in-house datasets and one publicly available dataset have demonstrated that AM-GAN consistently improves signal quality, significantly enhancing the accuracy of physiological parameter estimation both within-dataset and across datasets. Comparative analyses with other recent PPG enhancement methods, including $Err1(Err2)$, have confirmed the superior performance of AM-GAN in accurately calculating HR, RR, and SpO2 by producing clearer, more reliable pulsatile waveforms.
Moreover, AM-GAN has shown the ability to learn MA removal algorithms, specifically outperforming the original MR algorithm (MR algorithm, [b21zheng2023rapid]), particularly in scenarios involving complex interactions between cardiac and motion frequencies during intense physical activity. Although the current findings represent preliminary evidence, they clearly indicate substantial potential for future advancements.
Looking ahead, the AM-GAN framework could be expanded to integrate additional physiological signals, further enhancing its robustness and generalizability in various clinical and nonclinical scenarios. Future work could focus on applying this methodology to wearable and remote health monitoring systems, enabling real-time and accurate physiological assessments in environments with high MAs. Additionally, integrating AM-GAN with advanced analytics and deep learning techniques (such as self-supervised learning) could facilitate predictive monitoring, personalised healthcare solutions, and broader telehealth applications, thereby significantly extending the utility and impact of opto-physiological monitoring technologies.
Author contributions
XZ carried out the AM-GAN work with the associated comparison study and organized the manuscript. SH supervised, organized, and reviewed the study. VMD extracted multichannel signals and postprocessed the raw datasets. LB supervised the study of physiological monitoring and reviewed the manuscript. MD supervised the study of AM-GAN.
@@ Acknowledgments
The authors would like to acknowledge the support of Loughborough University in the conduct of the Ph.D. study during the period of October 2020 to December 2024. The authors also acknowledge a) the permit of the C2 dataset for SpO2 calculation from Carelight Limited, UK, b) the implementation of altitude study (Dec 2023 - Jan 2024) to obtain C2 dataset, led by Caroline Sunderland and James Donaldson in Nottingham Trent University, UK, and the shared datasets in the IEEE-SPC and PPGDalia.
@@
Appendix.
Table 1: Summary of prior MA removal methods.
Category Method Brief Summary Inputs Outputs Exercise Intensity Traditional / Signal-processing approaches Traditional EMD/ FFT/ ICA/ Wavelet [b4pankaj2022review] Classical decomposition / transform / denoising pipelines for MA reduction. PPG MA-reduced PPG; HR/ RR/ SpO2 Rest–low Traditional ANC(RLS) & AS-LMS [b5ram2011novel] Reference-based adaptive cancellation (RLS); variable step-size LMS. PPG++ACC MA-reduced PPG; HR/SpO2 General activity Traditional EPDA [b36bradley2024opening] Efficient envelope-based PPG denoising algorithm. PPG MA-reduced PPG; HR N/A Traditional TROIKA [b6zhang2014troika] Sparse spectrum reconstruction++spectral peak tracking with ACC assistance. PPG++ACC HR Low–high Traditional JOSS [b25zhang2015photoplethysmography] Jointly estimates PPG and accelerometer spectra via MMV-based sparse recovery. PPG++ACC HR Low–high Traditional CEEMDAN [b35zhang2024removal] CEEMDAN-based entropy feature and adaptive filter for removal. PPG++ACC Improved SpO2 accuracy; MA-reduced PPG Everyday activity Traditional ANFA [b18zheng2022adaptive] Adaptive frequency-domain MA removal with ACC assistance. PPG++ACC MA-free PPG; HR/RR Low–high Traditional MR [b21zheng2023rapid] Suppresses MAs by using mOEPS and ACC estimates to select a pulsatile component with minimal residuals. PPG++ACC ++VEL MA-free PPG; HR/RR/SpO2 Low–high Machine-learning / Deep-learning approaches ML/DL (End-to-end) End-to-end mapping [b8luque2018end] DL regression from raw PPG to target vital signs. PPG HR/RR/SpO2 N/A ML/DL (Features) Feature-extraction DL [b9biswas2019cornet] CNN-based feature learning from pre-processed PPG. PPG HR/RR/SpO2 N/A DL (End-to-end) DeepPPG [b29reiss2019deep] End-to-end model uses synchronized PPG/ACC spectrograms to estimate HR. PPG++ACC HR N/A DL (Quality) Robust 1D-CNN [b11goh2020robust] 1D-CNN signal-quality classifier (clean vs artefact). PPG Quality label; HR Low DL (Quality, 2D) 2D representations [b12chen2021signal] 2D transform ++ 2D-CNN for quality at rest/low intensity. 2D-PPG Quality label; HR Rest–low GAN (denoising) Corr-image GAN [b13afandizadeh2023accurate] Convert PPG to 2D correlation images; image GAN for MA removal. 2D-PPG MA-reduced PPG; HR Low GAN (PPG MA removal) PPG-GAN [b29zheng2022ppg] GAN to generate de-noised PPG without ACC/Gyro. PPG MA-reduced PPG; HR/RR Low–high Attention (HR) PULSE [b37kasnesis2023feature] Temporal convolutions ++ multi-head attention for HR. PPG HR N/A Attention (HR) Adaptive Q-PPG [b38kechris2024kid] Attention-enhanced Q-PPG for HR tracking. PPG HR N/A GAN++Attention (synthesis) CardioGAN [b31sarkar2021cardiogan] Attention generator; dual discriminators (time++freq) for ECG-from-PPG synthesis. PPG Synthesised ECG Rest GAN++Attention (EEG) EEGANet [b17sawangjai2021eeganet] GAN-based ocular artefact removal (EEG). EEG MA-reduced EEG N/A GAN++Attention (PPG) Attention GAN [b33sawangjai2025removal] Attention GAN with dual discriminator for MA removal without ACC/Gyro. PPG MA-reduced PPG; HR Low–medium This work (proposed) GAN++Attention (fusion) AM-GAN Attention within GAN for multi-sensor fusion and MA removal. PPG++ACC ++VEL MA-free PPG; HR/RR/SpO2 Low–high